OpenStudy (anonymous):

Use the Laws of Logarithms to combine the expression. log 54+1/2 log 6-log 6

OpenStudy (anonymous):

log54+log6^(1/2)-log6....... log[{(54+root(6))}/6)].....

OpenStudy (anonymous):

Should be \[log(\frac{54\sqrt{6}}{6})\] When simplified: \[log(9\sqrt{6})\] Rules of logarithms: \[log_b(m)+log_b(n)=log_b(mn)\] \[log_b(m)-log_b(n)=log_b(\frac{m}{n})\] And though it doesn't apply here: \[log_b(m^n)=nlog_b(m)\]